An iron bolt is used to connect two iron plates together. The bolt must withstand shear force up to about 3200N. Calculate the minimum diameter for the bolt on a safety factor of 8.0. (Shear strength of iron is 1.70x10^8 N/m^2

(Please WRITE out all ALL algebraic steps especially concerning finding the area of the bold. Please also show which numerical value should go in place of the diameter. Again, please do not skip any steps or do any mental math. I am having difficulties with the algebra.)
Answer #1

Total load, W = 3200 x 8 = 25600 N

Area, A = 25600/(1.7 x 10^8) = 1.51 x 10^-4 m^2

Diameter, d = sqrt(4A/pi) = sqrt(4 x 1.51 x 10^-4/pi) = 0.0138 m = 1.38 cm

Comment in case any doubt please rate my answer....